OpenStudy (anonymous):

Simple question about equation:

OpenStudy (anonymous):

\[z3=\frac{ -2-2j+j ^{2} }{ 2+2j-j ^{2} }\]

OpenStudy (anonymous):

What I must do here?

OpenStudy (perl):

is j the imaginary number , j^2=-1 ?

OpenStudy (anonymous):

yes

OpenStudy (perl):

go ahead and simplify , replace j^2 with -1

OpenStudy (anonymous):

Ok but I have the same difficult with x when I havent j and I HAVE X

OpenStudy (anonymous):

Ok I change it

OpenStudy (anonymous):

\[z3=\frac{ -2-2x+x ^{2} }{ 2+2x-x ^{2} }\]

OpenStudy (anonymous):

with this thing I have problems

OpenStudy (perl):

z_3 = (-2 -2j + j^2) / ( 2 + 2j -j^2) = ( -2 -2j -1 ) / ( 2 + 2j - (-1) ) = ( -3 -2j) / ( 3 +2j) = -(3 + 2j) / (3 + 2j) = -1

OpenStudy (anonymous):

you multiplicate with -1?

OpenStudy (perl):

with x you can try factoring it

OpenStudy (perl):

I factored out -1 from (-3 -2j ) (-3 -2j ) = - ( 3 + 2j)

OpenStudy (anonymous):

Ok I cant understan how function this factor

OpenStudy (perl):

The numerator is given as (-2 -2x +x^2) but -1( 2 +2x -x^2) = -2 -2x + x^2 check it for yourself, distribute

OpenStudy (anonymous):

a ok I understand and with x?

OpenStudy (anonymous):

I dont understand because I cant simplyfi with x

OpenStudy (anonymous):

Hi

OpenStudy (perl):

what do you mean simplify? you can factor out -1 , and then cancel

OpenStudy (perl):

you mean factor?

OpenStudy (perl):

it is not necessary to factor the quadratic

OpenStudy (anonymous):

\[z3=\frac{ -2-2x+x ^{2} }{ 2+2x-x ^{2} }\]

OpenStudy (perl):

replace the numerator -2 -2x + x^2 with -( 2 + 2x -x^2)

OpenStudy (anonymous):

when I have this I cant simplyfi

OpenStudy (anonymous):

I cant understand why I cant simplify +x^2 with x2 and other number

OpenStudy (anonymous):

x^2/-x^2

OpenStudy (perl):

do you agree that -2 -2x + x^2 = - ( 2 + 2x -x^2)

OpenStudy (perl):

because of order of operations. you cannot divide before adding all the terms in the numerator

OpenStudy (anonymous):

ok and wher I can find other exercis of this? what I must search on google beacuse I have difficulty

OpenStudy (anonymous):

\[z3=\frac{ -2-2x+x ^{2} }{ 2+2x-x ^{2} } ~~ I ~~ mean ~~ other ~~ exercise~~ like ~~ this\]

OpenStudy (perl):

can you finish the problem simplifying